OpenStudy (anonymous):

Determine whether the graphs of the following equations are parallel or perpendicular? 2x-8y= -24, 4x + y= -2, x-4y= 4

OpenStudy (anonymous):

Parallel lines have equal slopes. Perpendicular lines have slopes that are the negatice reciprocal of each other. Take each line and compute its slope. Then use the principles outlined above.

OpenStudy (anonymous):

I am guessing they a parallel

OpenStudy (anonymous):

Don't guess...work out the slope for each. You can show all your work and I can take a look at it.

OpenStudy (anonymous):

I graphed it and the 1st one and third one are parallel

OpenStudy (anonymous):

Sometimes graphs can be deceiving...best way to go is to take each line and write it in the form of y = mx + b, and m = slope. I'll do one of the lines: x - 4y = 4 We want to solve for y: subtract x from both sides, you get - 4y = 4 -x Now divide both sides by - 4, you get y = (4 - x)/-4 So the slope is the coefficient of x, and that is m. coefficient of x is -1/-4 = 1/4. So this line had slope of 1/4 Now take the

OpenStudy (anonymous):

Now you find the slope of the first line.

OpenStudy (anonymous):

Would x equal -12 + 8y

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

you want to solve for y.

OpenStudy (anonymous):

Y=3- 2x

OpenStudy (anonymous):

Can you just tell me I need to finish this today